OpenStudy (anonymous):

Suppose that f(0)=3 and f′(x)≤4 for all values of x. Use the Mean Value Theorem to determine how large f(4) can possibly be.

myininaya (myininaya):

What happened when you tried to apply the M.V.T. ?

OpenStudy (perl):

by MVT we have [f(4) - f(0)]/ ( 4 - 0) = f' (c)

jimthompson5910 (jim_thompson5910):

Let f ' (x) be the largest it can be. So let f ' (c) = 4 for some value of c Also, let f(4) = x By the mean value theorem, we know \[\Large \frac{f(4)-f(0)}{4-0} = f \ '(c)\] \[\Large \frac{x-3}{4-0} = 4\] solve for x

OpenStudy (perl):

f(4) - 3 = f ' (c) * 4 f(4) = f ' (c) *4 + 3 now you are given that f ' (x) <= 4 for all x. this means f(4) <= 4 * 4 + 3 , by substitution